# YITP-24-51

### v1

@@ 1 Introduction
Of course, we can argue that the $\epsilon$ expansion is just a mathematical trick and we should not speak of its physical meaning. There are infinitely different ways to approach the Heisenberg fixed point by changing the dimensionality, and in a certain sense, there is no “physically more reasonable” model. Still, it seems an interesting question to ask what happens if we study the $O(4-\epsilon)$ two-form (or anti-symmetric matrix) model in $4-\epsilon$ dimensions rather than the $O(3)$ vector model? It may give an intrinsic ambiguity in the predictions of the $\epsilon$ expansions by taking the $\epsilon\to 1$ limit. A related question has been addressed when we study (space-time) “spinors” in $4-\epsilon$ dimensions. The spinor representation in non-integer dimensions is not well-defined and many different prescriptions exist. The role of so-called evanescent operators there has not been completely understood (see e.g. [2]).
@@ 3.3 More bootstrapping
Figure 7: Conformal bootstrap bound with $O(3)$ symmetry in $d=3.98$. Black $S$. Red $A_{4}$. Blue $T$. Orange $B_{4}$. The horizontal axis is the conformal dimension of a spin-zero operator in the anti-symmetric representation. The bounds in the $T$ and $S$ sectors are the same as those of the vector model and the kinks are located at the $O(3)$ vector model $V$ represented by Circle. We do not know a candidate for the kinks in the $A_{4}$ and $B_{4}$ sectors. The triangles are predictions of evanescent operators from $M_{-}$. The predictions from $M_{+}$ are not shown here, which are located at $\Delta_{\phi}=0.99$ and violate the bound.
@@
Next, let us discuss the $O(6)$ case in $d=3.98$ presented in fig 8. The $\epsilon$ expansion of the anti-symmetric matrix model shows that the matrix model fixed points $M_{\pm}$ are complex fixed points and the only real fixed point is the $O(15)$ vector model. We observe that the kink in the $B_{4}$ sector is precisely saturated by the $T$ operator of the $O(15)$ vector model. It is uncommon to observe an enhancement of the symmetry in the bound of the non-singlet operator. The fact that the fixed point that saturates the $B_{4}$ sector is switched from $M_{-}$ (when $N\sim 4$) to $V$ (when $N=6$) should coincide with $6>N_{c}$ when $\epsilon$ is small.
@@
We finally present physically interesting bound with $O(4)$, $O(5)$ and $O(6)$ symmetries in $d=3$. See Fig 9, 10 and 11. They could be realized in nature as the Landau-Ginzburg model with $O(N)$ anti-symmetric matrix order parameter. In the $O(4)$ case, we find that the $T$ operator in the vector model fixed point with the enhanced $O(6)$ symmetry does not saturate any bound. The kink in the $A_{4}$ sector around $\Delta_{\phi}=0.516$ and another one in the $B_{4}$ sector around $\Delta_{\phi}=0.519$ may indicate a nontrivial matrix model fixed point such as $M_{-}$ or potentially $M_{+}$. Compare them with the $O(3.98)\sim O(4)$ in $d=3.98$ in Fig 2, where $M_{-}$ seems to be located at kinks of $A_{4}$ and $B_{4}$. There, we have also seen that the $T$ operator in the vector model fixed point does not saturate any bound.
@@
$d$ $3$ $3.98$ $3$ $3.98$ $3$ $3$ $3.98$
$N$ $3$ $3$ $4$ $3.98$ $5$ $6$ $6$
$S$ $V$ $V$ $V$ $V$ $V$ $V$ $V$
$T$ $V$ $V$ very weak ? very weak very weak ?
$A_{4}$ $M-$? weak weak ?
$B_{4}$ $M_{-}$($M_{+}$?) $V$? $V$ $V$

### v2

@@ 1 Introduction
Of course, we can argue that the $\epsilon$ expansion is just a mathematical trick and we should not speak of its physical meaning. There are infinitely different ways to approach the Heisenberg fixed point by changing the dimensionality, and in a certain sense, there is no “physically more reasonable” model. Still, it seems an interesting question to ask what happens if we study the $O(4-\epsilon)$ two-form (or anti-symmetric matrix) model in $4-\epsilon$ dimensions rather than the $O(3)$ vector model? It may give an intrinsic ambiguity in the predictions of the $\epsilon$ expansions by taking the $\epsilon\to 1$ limit. A related question has been addressed when we study (space-time) “spinors” in $4-\epsilon$ dimensions. The spinor representation in non-integer dimensions is not well-defined and many different prescriptions exist. The role of so-called evanescent operators there has not been completely understood (see e.g. [2]). See also [3] for the evanescent operators with $O(N)$ symmetry, which are relevant to our study.
@@ 3.3 More bootstrapping
Figure 7: Conformal bootstrap bound with $O(3)$ symmetry in $d=3.98$. Black $S$. Red $A_{4}$. Blue $T$. Orange $B_{4}$. The horizontal axis is the conformal dimension of a spin-zero operator in the anti-symmetric representation. The bounds in the $T$ and $S$ sectors are the same as those of the vector model and the kinks are located at the $O(3)$ vector model $V$ represented by Circle. We do not know a candidate for the kinks in the $A_{4}$ and $B_{4}$ sectors. The triangles are predictions of evanescent operators from $M_{+}$. The predictions from $M_{-}$ are not shown here, which are located at $\Delta_{\phi}=0.99$ and violate the bound.
@@
Next, let us discuss the $O(5)$ and $O(6)$ cases in $d=3.98$ presented in fig 8 and fig 9. The $\epsilon$ expansion of the anti-symmetric matrix model shows that the matrix model fixed points $M_{\pm}$ are complex fixed points and the only real fixed point is the $O(15)$ vector model. We observe that the kink in the $B_{4}$ sector is precisely saturated by the $T$ operator of the $O(10)$ and $O(15)$ vector model. It is uncommon to observe an enhancement of the symmetry in the bound of the non-singlet operator. The fact that the fixed point that saturates the $B_{4}$ sector is switched from $M_{-}$ (when $N\sim 4$) to $V$ (when $N=5,6$) should coincide with $5>N_{c}$ when $\epsilon$ is small.
@@
We finally present physically interesting bound with $O(4)$, $O(5)$ and $O(6)$ symmetries in $d=3$. See Fig 10, 11 and 12. They could be realized in nature as the Landau-Ginzburg model with $O(N)$ anti-symmetric matrix order parameter. In the $O(4)$ case, we find that the $T$ operator in the vector model fixed point with the enhanced $O(6)$ symmetry does not saturate any bound. The kink in the $A_{4}$ sector around $\Delta_{\phi}=0.516$ and another one in the $B_{4}$ sector around $\Delta_{\phi}=0.519$ may indicate a nontrivial matrix model fixed point such as $M_{-}$ or potentially $M_{+}$. Compare them with the $O(3.98)\sim O(4)$ in $d=3.98$ in Fig 2, where $M_{-}$ seems to be located at kinks of $A_{4}$ and $B_{4}$. There, we have also seen that the $T$ operator in the vector model fixed point does not saturate any bound.
@@
$d$ $3$ $3.98$ $3$ $3.98$ $3$ $3.98$ $3$ $3.98$
$N$ $3$ $3$ $4$ $3.98$ $5$ $5$ $6$ $6$
$S$ $V$ $V$ $V$ $V$ $V$ $V$ $V$ $V$
$T$ $V$ $V$ very weak ? very weak ? very weak ?
$A_{4}$ $M-$? weak ? weak ?
$B_{4}$ $M_{-}$($M_{+}$?) $V$? $V$ $V$ $V$